Mass of BaSO4 is 0.075g which I calculated to be 3.21x10^-4 moles of BaSO4

I want to find the Moles, Mass and Percent by mass of the sulfate ION?
Answer #1

1 mole of BaSO4 contains 1 mole of Ba2+ & 1 mole of SO42- ions

So 3.21*10-4 moles of BaSO4 contains 3.21*10-4 moles of SO42- ion

Molar mass of BaSO4 = atomic mass of Ba + Atomic mass of S + (4* atomic mass of O)

                                 = 137.3 + 32 +(4*16) g/mole

                                = 233.3 g/mole

Molar mass of SO42- ion = atomic mass of S + ( 4*atomic mass of O)

                                     = 32+(4*16)

                                    = 96 g/mole

1 mole = 233.3 g of BaSO4 contains 96 g of sulfate ion

0.075 g of BaSO4 contains (0.075g*96g)/233.3 g = 0.031 g of sulfate ion

Mass percent of sulfate ion = ( mass of sulfate ion / mass ofBaSO4) *100

                                       = (0.031 g / 0.075g) *100

                                      = 41.1%